OpenStudy (anonymous):

evaluate the integral using the given substitution. x^4(x^5-3)^5 dx, u=x^5-3 I got: 1/25(x^5-3)^5+C this is correct?

OpenStudy (mimi_x3):

I don't think that the answer you got is right.. I think you made amistake here: \[\frac{1}{5} \int\limits u^{5}du \] when you integrate it..

OpenStudy (anonymous):

1/5(x^5-3)^5+C ??

OpenStudy (mimi_x3):

Nope.. use: \[\int\limits x^{a} dx = \frac{x^{a+1}}{a+1} +C\]

OpenStudy (anonymous):

\[\frac{1}{5}\int u^5du\] \[\frac{1}{30}u^6\]

OpenStudy (anonymous):

hi mimi nice bike!

OpenStudy (mimi_x3):

lol, people are starting the bike trend now :P

OpenStudy (anonymous):

wanna race?

OpenStudy (mimi_x3):

haha, I will lose for sure!

OpenStudy (anonymous):

i dunno your bike looks newer than mine

OpenStudy (anonymous):

those are my two choices, and these two. 1/25x^25-3+C 1/15(x^5-3)^3+C

OpenStudy (mimi_x3):

my books looks new but your one looks nicer! Are you sure about the choices?

OpenStudy (anonymous):

unfortunately...yes i'm sure :/

OpenStudy (anonymous):

i get \[\frac{1}{30}(x^5-3)^6\]

OpenStudy (mimi_x3):

I get that as well; dont forget about +C

OpenStudy (anonymous):

darn...that's not even one of my choices lol

OpenStudy (anonymous):

this is the integral \[\int x^4(x^5-3)^5dx\] right?

Parth (parthkohli):

Hey satellite, Imma bike too :)

OpenStudy (anonymous):

wow lots of bikes today one day everyone was my bike

OpenStudy (anonymous):

yes, using the substitution u=x^5-3 all these cool bike peeps helping me are so cool. thank you! :)

Parth (parthkohli):

@satellite73 You are our idol :) you don't even know me I guess. I am Parth.

OpenStudy (anonymous):

Penny-Farthing ftw